the length of a rectangle is 10 inches more than the width. the perimeter of the rectangle is 40 inches. find the width of the rectangle

Answer by skthakur(2)

let width = X
so lenght X+10
Now
2(X+X+10)=40
4X+20=40
4X=40-20
X=20/4
X=5
Width=5 inches